For each situation, present the accounts and dollar amounts that would appear on comparative balance sheets and income statements for the ears ending 12/31/16 and 12/31/15.

Prepare an amortization table

List all accounts and dollar amounts. Round dollar amounts to the nearest dollar. You do not need to include cash.

For the Classification {Class:} column of the Balance Sheet use:

• A for Asset

• L for Liability

• E for Equity

For the Classification {Class:} column of the Income Statement use:

• R for Revenue

• E for Expense

The following information relates to R-U Ready Company, a publicly traded company:

  1. R. U. Ready acquired Machine B on 1/1/13 by signing a 9% installment note to be paid in 5 equal installments of $65,000. Each payment is due on 12/31, with the first payment due 12/31/13. The useful life of Machine B is 8 years with no residual value. Assume double declining balance depreciation.
  2. On 1/1/10, R-U Ready issued $100,000, 6.5%, 10-year bonds at an effective rate of 4.75%. Interest is paid annually on 12/31 of each year.
  3. R-U Ready acquired land with a FMV of $220,000 on 1/1/14 by issuing a 3-year non-interest-bearing note due 12/13/16. Assume 8% as the effective rate of interest. (You are not provided the face value of the note, but have enough information to determine it.)

Please answer question number 2 and 3.
Answer #1
Case-1
The cost or present value of the machine can be calculated uisng the PV of ordinary annuity formula,
where, the equal annuities = 65000
at the interest rate r=9%
& no.of payments= 5
So, the cost of the machine is:
Cost/PV=65000*(1-1.09^-5)/0.09=
252827
The amortisation table for the above Installment Note will be:
Date Annuity Tow.Int. Tow.Prin. Prin. Bal.
1 2 3=Prev.5*9% 4=2-3 5=Prev.5-4
1/1/2013 252827
12/31/2013 65000 22754 42246 210581
12/31/2014 65000 18952 46048 164534
12/31/2015 65000 14808 50192 114342
12/31/2016 65000 10291 54709 59633
12/31/2017 65000 5367 59633 -1
Total 325000 72172 252828
Depreciation under DDB method
Depreciable value=the carrying value
Depreciation under DDB =2*St.line rate of depn.
ie. 2*(1/8 yrs.)= 25%
So, drawing up the annual depn. schedule
for the 8 yr. useful life of the machine,
Date/Yr. ending Depn. For the Year(25%*Previous carrying value) Carrying Value
1/1/2013 252827
12/31/2013 63207 189620
12/31/2014 47405 142215
12/31/2015 35554 106661
12/31/2016 26665 79996
12/31/2017 19999 59997
12/31/2018 14999 44998
12/31/2019 11249 33748
12/31/2020 8437 25311
Income Statement(Partial)
12/31/2016 12/31/2015
(E)Depreciation 26665 35554
(E)Interest on Installment note 10291 14808
Balance Sheet(Partial)
Liabilities 12/31/2016 12/31/2015 Assets 12/31/2016 12/31/2015
(L)9% Installment Note 59633 114342 (A) Machinery-Gross 252827 252827
Less: Acc. Depn. 172831 146166
Machinery, Net 79996 106661
(E)Shareholders' Equity -36956 -50362
(Depn.+Int.)
Case-2
The PV of the bond/issue price of the bond Is calculated using the PV of annuity formula as above,
where,
the Face value of the bond= 100000
Annual coupon amt.=100000* 6.5%=6500
Effective rate of interest = 4.75% p.a.
no.of coupon payments= 10
so, the PV or issue value of the bond=
Pv of the bond=PV of all its future coupons+PV of fcae value to be received at maturity
(6500*(1-1.0475^-10)/0.0475)+(100000/1.0475^10)=
113679
So, the bond is issued at a premium of 113679-100000=13679
Now the effective interest rate amortisation table of the bond over the 10 yrs.Will be
A B C D E F G
Date Coupon int.(6.5%*FV) Interest expense(4.75%*Prev. BV in G) Premium amortised(C-B) Bal.in Bond Premium a/c Bond payable BV of bonds(F+E)
Cr. Cash Dr. Int. exp. Dr. Bond Premium
1/1/2010 13679 100000 113679
12/31/2011 6500 5400 -1100 12579 100000 112579
12/31/2012 6500 5347 -1153 11426 100000 111426
12/31/2013 6500 5293 -1207 10219 100000 110219
12/31/2014 6500 5235 -1265 8954 100000 108954
12/31/2015 6500 5175 -1325 7630 100000 107630
12/31/2016 6500 5112 -1388 6242 100000 106242
12/31/2017 6500 5047 -1453 4789 100000 104789
12/31/2018 6500 4977 -1523 3266 100000 103266
12/31/2019 6500 4905 -1595 1671 100000 101671
12/31/2020 6500 4829 -1671 1 100000 100001
Income Statement(Partial)
12/31/2016 12/31/2015
(E)Interest expense 5112 5175
Balance Sheet(Partial)
12/31/2016 12/31/2015
(L)6.5% 10-Yr. Bonds payable 100000 100000
Unamortised Bond premium 6242 7630
106242 107630
(E)Shareholders' Equity -5112 -5175
(Int.exp.)
Case-3
The Face Value of the note =220000*1.08^3=
277137
Date Interest exp./ payable Total balance
1/1/2014 220000
12/31/2014 17600 237600
12/31/2015 19008 256608
12/31/2016 20529 277137
Income Statement(Partial)
12/31/2016 12/31/2015
(E)Interest expense 20529 19008
Balance Sheet(Partial)
Liabilities 12/31/2016 12/31/2015 Assets 12/31/2016 12/31/2015
(L)3-Yr. Note 220000 Land 220000 220000
(L)Interest payable 36608
0 256608
(E)Shareholders' Equity -20529 -19008
(Int.exp.)